OpenStudy (anonymous):

Find the slop of a urve y=(x^3)-2x-2 at a point p(2,-2). How do I solve this?

OpenStudy (anonymous):

Just Differentiate it and then put the points

OpenStudy (anonymous):

meant curve, not urve

OpenStudy (anonymous):

\[y = x^3 -3x -2 \] \[\frac{dy}{dx} = 3x^2 - 2 \]

OpenStudy (anonymous):

Now put the point and you will have the slope

OpenStudy (anonymous):

I tried that and thought it was 0. But thats not the answer

OpenStudy (anonymous):

\[y' = 3( 2)^2 -2 \] \[y' = 12-2\] \[y' =10\]

OpenStudy (anonymous):

hmm how'd you get zero?

OpenStudy (anonymous):

If \(f(x)\) is a function or some curve then \(f'(x)\) is it's first derivative and \(f'(a)\) is slope of tangetn at point where x co-ordinate is a.

OpenStudy (anonymous):

Well my equation isnt the same, I just was trying to figure out how to do it. The real one is y=(x^2)-2x-2. I got somehow 0

OpenStudy (anonymous):

tangent *

OpenStudy (anonymous):

hmm \[y = x^2 - 2 -2 \]

OpenStudy (anonymous):

Just was trying to figure how out how its done. Still nothing. I did a formula you did. Got 6. But thats wrong

OpenStudy (anonymous):

\[\frac{dy}{dx} = 2x -2 \] \[ \frac{dy}{dx} = 4-2=2\]

OpenStudy (anonymous):

I will try that. Honestly wish how it works though.. Cant figure it out

OpenStudy (anonymous):

I will try that. Honestly wish how it works though.. Cant figure it out

OpenStudy (anonymous):

OKay

OpenStudy (anonymous):

hmm what's the exact question curve is \[y = x^2 -2x-2\] right? so that makes \[y' =2x -2 \] now what are the points ?

OpenStudy (anonymous):

You got it right with 2. but why? im more worried about the reason. Yes the curve is right

OpenStudy (anonymous):

hmm reason you can watch mit ocw lecture 1 for that a very clear explanation

OpenStudy (anonymous):

single variable calculus

OpenStudy (anonymous):

Thanks I will keep just checking my book

OpenStudy (anonymous):

Good Luck : )

OpenStudy (anonymous):

So wait how do I find the equation of a tangent line?

OpenStudy (anonymous):

hmm for finding that hmm you have slope right ! so the equation becomes y=mx+c now in this case slope is 2 y = 2x +c we also know that line passes through (2,-2) so just put in the points in the equation to determine c . -2 = 2x2 + c c = -6 y=2x - 6 you have your equation

OpenStudy (anonymous):

damn.. I need to figure out how this works. Just dont see it yet